The present value of a loan in which ​$3000 is to be paid out a year from today with the interest rate equal to 22​% is _________?

​(Round your response to the neareast two decimal​ place)

If a loan is paid after two​ years, and the amount ​$5000 is to be paid then with a corresponding 22​% interest​ rate, the present value of the loan is ______?

​(Round your response to the neareast two decimal​ place)

______________ is based on the notion that a dollar paid in the future is less valuable than a dollar paid today.

A) Cash Flow

B) Present Discount Value

C) Interest Rate
Answer #1

Ans: The present value of a loan in which ​$3000 is to be paid out a year from today with the interest rate equal to 22​% is $2459.1

Explanation:

Future value ( F ) = $3000

Number of interest period ( n ) =1

Interest rate = 22%

Present value = F ( P / F , i ,n )

=$3000 ( P/F , 22% , 1 )

= $3000 * 0.8197

= $2459.1

Or

P = F / ( 1 + i )n

= $3000 / ( 1+ 0.22 )1

= $3000 / (1.22)1

= $3000 / 1.22 = $2459

Ans: If a loan is paid after two​ years, and the amount ​$5000 is to be paid then with a corresponding 22​% interest​ rate, the present value of the loan is $3359.5

Explanation:

Future value ( F ) = $5000

Number of interest period ( n ) =2

Interest rate = 22%

Present value = F ( P / F , i ,n )

=$5000 ( P/F , 22% , 2 )

= $5000 * 0.6719

= $3359.5

Ans: Present Discount Value

Explanation:

Present Discount Value is based on the notion that a dollar paid in the future is less valuable than a dollar paid today.